# Conformal Temporal Logic Planning using Large Language Models

### v4

@@
Abstract
This paper addresses planning problems for mobile robots. We consider missions that require accomplishing multiple high-level sub-tasks, expressed in natural language (NL), in a temporal and logical order. To formally define the mission, we treat these sub-tasks as atomic predicates in a Linear Temporal Logic (LTL) formula. We refer to this task specification framework as LTL-NL. Our goal is to design plans, defined as sequences of robot actions, accomplishing LTL-NL tasks. This action planning problem cannot be solved directly by existing LTL planners because of the NL nature of atomic predicates. To address it, we propose HERACLEs, a hierarchical neuro-symbolic planner that relies on a novel integration of (i) existing symbolic planners generating high-level task plans determining the order at which the NL sub-tasks should be accomplished; (ii) pre-trained Large Language Models (LLMs) to design sequences of robot actions based on these task plans; and (iii) conformal prediction acting as a formal interface between (i) and (ii) and managing uncertainties due to LLM imperfections. We show, both theoretically and empirically, that HERACLEs can achieve user-defined mission success rates. Finally, we provide comparative experiments demonstrating that HERACLEs outperforms LLM-based planners that require the mission to be defined solely using NL. Additionally, we present examples demonstrating that our approach enhances user-friendliness compared to conventional symbolic approaches.
Index Terms:
@@ Abstract
I Introduction
@@
Designing autonomous agents with task planning capabilities is a long standing goal in robotics and automation [1, 2, 3, 4]. Achieving this goal requires the development of mission specification frameworks and planning algorithms capable of generating plans—sequences of robot actions—to accomplish assigned missions. These frameworks should allow practitioners to define missions unambiguously and in a user-friendly manner, while planners need to exhibit computational efficiency and be supported by correctness guarantees.
Recently, several task planners have been proposed that can design correct-by-construction plans for complex missions with specified temporal and logical requirements using Linear Temporal Logic (LTL) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, this framework demands a significant amount of expertise and manual effort to rigorously define complex tasks. This is because defining LTL tasks requires specifying multiple atomic predicates (i.e., Boolean variables) to model desired low-level robot configurations and coupling them using temporal/Boolean operators. Additionally, the more complex the task requirements, the larger the number of predicates and temporal/logical operators needed to define the corresponding LTL formula. This not only compromises the user-friendliness of the framework but also increases the computational cost of designing robot plans [14].
@@
On the other hand, Natural Language (NL) has emerged as a more user-friendly method for specifying robot missions. Early research in NL-based planning primarily focused on mapping NL to planning primitives [30, 31, 32, 33, 34] using statistical machine translation [35] to identify data-driven patterns for translating free-form commands into a formal language defined by a grammar. However, these approaches were limited to structured state spaces and simple NL commands. Motivated by the remarkable generalization abilities of pre-trained Large Language Models (LLMs) across diverse task domains [36, 37, 38, 39], there has been increasing attention on utilizing LLMs for NL-based planning [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. While LLM-based planning methods are more user-friendly than LTL-based planners, they have two main limitations: (i) they lack mission performance guarantees; and (ii) their ability to design correct plans deteriorates with increasing task complexity [57].
@@
In this paper, we propose an alternative approach to mission specification and robot planning aimed at addressing some of the limitations observed in current LTL and NL planning frameworks. Our approach involves tightly coupling LTL and NL to leverage their respective strengths. First, we discuss our proposed mission specification framework. We model a complex task as comprised of a collection of sub-tasks, with temporal and logical relationships among them represented using an LTL specification. The sub-tasks thus serve as atomic predicates in an LTL formula, with the associated predicate being true if the sub-task is successfully accomplished, and false otherwise. Our key departure from conventional LTL specifications is that we specify each such sub-task in NL, rather than directly grounding it in the low-level configuration space; see e.g., [10, 11, 14, 15, 24]. For example, consider a simple task in which the robot first needs to deliver a bottle of water to the kitchen table, and only afterwards deliver a coke to the office desk. We can define two NL sub-tasks, ‘deliver a bottle of water to the kitchen table’ and ‘deliver a coke to the office desk’, and use an LTL formula to express the logical and temporal relationship between them. The advantage of this framework over conventional LTL is that LTL-NL formulas require fewer predicates and temporal/logical operators while capturing the same task, making them more user-friendly to define. In comparison to NL instructions (which are more user-friendly to define than LTL-NL commands), our framework enables the automatic decomposition of the overall mission (i.e., an LTL-NL specification) into multiple NL-based sub-tasks using existing symbolic planners [19, 20, 26]. It is important to note that such a decomposition is highly challenging to perform in a correct-by-construction fashion for missions expressed exclusively using NL. As we will discuss later, task decomposition plays a critical role in enabling our planning framework to consistently outperform existing LLM-based planners, especially as mission complexity increases.
Figure 1: This paper proposes HERACLEs, a neuro-symbolic planning framework for mobile robots with LTL-NL missions. The framework starts with a symbolic task planner that generates the next language-based sub-task that should be accomplished to make mission progress. Pre-trained LLMs are then deployed to design robot plans for that sub-task. The LLM planner seeks assistance from the symbolic planner either when the robot plan cannot be executed due to environmental constraints or when it is uncertain about how to accomplish that sub-task. Uncertainty of the LLM is quantified using conformal prediction (CP). If the LTL task planner cannot assist, then help from users is requested (not shown). The mission in this figure requires the robot to eventually deliver a pen to location D but only after either an apple or a drink has been delivered to D. The symbolic planner first generates the sub-task ‘deliver a drink to E’. Since the prediction set using CP is non-singleton (including the actions ‘Go to location A’ and ‘Go to location D’), the LLM-based planner asks the symbolic planner to generate an alternative sub-task to make progress. The new task is ‘deliver an apple to E’; see also Example II.4.
@@
Second, we address the challenge of designing robot plans that ensure correctness with respect to LTL-NL formulas. Specifically, we consider a mobile robot equipped with various skills (e.g., mobility, manipulation, and sensing) tasked with missions expressed as LTL-NL specifications. Each predicate in the LTL-NL formula is satisfied if an NL-based sub-task is accomplished, requiring the robot to apply its skills to various semantic objects and regions of interest in the environment. For instance, the sub-task ‘deliver a bottle of water to the kitchen table’ requires the robot to move to the location where the bottle is, grasp it, move to the kitchen table, and release the object. Observe that a predicate in an LTL-NL formula is satisfied by a sequence of robot configurations/actions. While we assume that the locations and semantic labels of these objects are known, the geometric environmental structure is initially unknown. The latter may hinder the completion of certain NL-based sub-tasks (e.g., when an object is inaccessible). Our goal is to design robot plans, defined as a sequence of robot actions, satisfying the assigned LTL-NL missions.
@@
To address this task planning problem, we propose a hierarchical neuro-symbolic planner that harnesses the complementary strengths of LLMs and symbolic planners. Specifically, our proposed planner leverages existing LTL planners that dynamically determine, based on the current mission status, the next NL-based sub-task the robot should accomplish to make progress in the mission. In our implementation, we select the symbolic planner proposed in [19, 20] due to its computational efficiency; however, we emphasize that any other symbolic planner can be used. Then, we employ pre-trained LLMs to generate robot plans satisfying the NL sub-task announced by the symbolic planner [37, 58]. If the geometric environmental structure prevents the robot from executing these plans (e.g., obstacles blocking the entrance to regions of interest), the LTL task planner will generate alternative sub-tasks (if they exist) as in [19, 20]. A challenge here is that LLMs tend to hallucinate, i.e., to confidently generate incorrect outputs [59]. Such false confidence in incorrect outputs poses a significant challenge in LLM-based robot planning. To reason about the correctness of the LLM-generated plans, inspired by [41, 60], we leverage conformal prediction (CP) [61, 62, 63], a statistical tool for uncertainty quantification in black-box models that has been recently applied to various safe autonomy tasks [64, 65, 60, 41, 66, 67, 68, 69, 70]. CP constructs on-the-fly prediction sets that contain the correct robot action with user-specified confidence. This allows the LLM to determine when it is uncertain about its predictions. In cases of high uncertainty, indicated by non-singleton prediction sets, the robot seeks assistance as follows. First, a help-message is sent to the symbolic planner requesting alternative sub-tasks to make mission progress. If such a sub-task exists, it is forwarded to the LLM planner. Otherwise, the robot asks for help from a user about how to proceed. This formal interface between the symbolic planner and the pre-trained LLM gives rise to our planner, called HERACLEs, for HiERArchical ConformaL natural languagE planner; see Fig. 1. The generated plans are executed using existing low-level controllers. We show both theoretically and empirically that HERACLEs can achieve user-specified theoretical mission success rates while requiring low help rates.
@@
Related works: (i) As discussed earlier, a substantial body of symbolic planners exists that can design plans satisfying LTL tasks [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. We emphasize that these planners cannot solve the considered LTL-NL planning problem due to the NL nature of the predicates. Specifically, these works can design a sequence of NL-based predicates that need to be satisfied to accomplish the LTL-NL mission; however, they cannot design a sequence of robot actions (e.g., ‘grab the bottle of water at location A’, ‘go to the kitchen table’, ‘drop off the bottle’) that satisfies a NL-based predicate (e.g., ‘deliver a bottle of water to the kitchen table’), let alone an LTL-NL formula, as they cannot process NL. We note again that our approach is tightly integrated with these symbolic planners as they are used to generate a sequence of NL sub-tasks that should be accomplished to satisfy an LTL-NL formula. (ii) Related are also recent neural planners that can design plans for tasks, described solely in NL, using pre-trained LLMs; see e.g., [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]; a recent survey can be found in [71, 72, 73]. However, the ability of LLMs to design correct plans drops significantly as the mission complexity increases as shown in [57, 74] as well as in our experiments. Additionally, as discussed earlier, the majority of these planners lack correctness and mission completion guarantees. Among these works, probably the closest one is [41] as it also applies CP for uncertainty alignment of LLMs allowing robots to ask for help from users in order to achieve desired mission completion rates. However, as shown in our comparative experiments, HERACLEs requires significantly lower help rates than [41] due to its task decomposition feature enabled by the LTL-NL specification. (iii) Neuro-symbolic planners that integrate LLMs with symbolic planners have also been proposed recently in [57, 75, 76, 77, 78, 79, 80]. Their key idea is to employ LLMs to translate NL commands into temporal logic specifications. Once this translation is completed, existing symbolic planners can be employed to design correct plans. A key limitation is that the translation process is not supported by correctness guarantees. As a result, the resulting plan may not satisfy the original NL instructions. In contrast, HERACLEs employs LLMs and symbolic planners in a fundamentally different way while CP allows the design of plans to achieve user-specified mission completion rates.
@@
Comparative Evaluations: We provide extensive comparative experiments showing that HERACLEs outperforms state-of-the-art LLM-based planners in terms of planning performance. In these experiments, we manually translate LTL-NL missions into NL instructions which serve as inputs to these planners. First, we compare HERACLEs against the LLM-based planners proposed in [40, 50, 81], which, however, do not utilize CP and do not allow robots to ask for help. Our comparative experiments show that the ability of these approaches to design correct plans tends to decrease as the number of temporal and logical requirements in the mission increases. To the contrary, the planning performance of HERACLEs consistently outperforms these baselines (even when the help mode is deactivated). Second, we compare our planner against a conformalized planner [41] that also allows robots to ask for help to achieve desired mission success rates. Our empirical analysis demonstrates that HERACLEs achieves significantly lower help rates than [41] to attain the same mission success rate. We attribute the performance gap between these baselines [40, 50, 81, 41] and HERACLEs primarily to the LTL-NL framework and the symbolic planner decomposing missions into multiple smaller/simpler NL-based sub-tasks that LLMs can handle more efficiently. Finally, we provide examples demonstrating the user-friendliness of LTL-NL formulas over the corresponding LTL specifications. Specifically, we show that LTL-NL formulas require significantly fewer predicates than an LTL formula modeling the same task. This difference in the ‘length’ of LTL and LTL-NL formulas becomes more pronounced as the complexity of the mission requirements increases.
@@
Contribution: The contribution of the paper can be summarized as follows. First, we propose a new task specification approach, called LTL-NL, to define complex high-level robot tasks. The advantage of LTL-NL over conventional LTL is its user-friendliness as it requires fewer predicates to define the same task while the predicates can be defined using NL, rather than low-level system configurations. Second, we introduce HERACLEs, a new neuro-symbolic planner that integrates existing symbolic planners with pre-trained LLMs to design plans satisfying missions encoded as LTL-NL formulas. Third, we show, both theoretically and empirically, that HERACLEs can achieve user-specified mission success rates through a formal interface, employing CP, between the LLM and the symbolic planner. Fourth, we provide extensive comparative experiments highlighting that HERACLEs outperforms state-of-the-art LLM-based planners in terms of its ability to design correct plans and achieve lower help rates.
II Problem Formulation
Robot System and Skills: Consider a robot governed by the following deterministic and known dynamics:
${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ (1)

### v5

@@
Abstract.
This paper addresses temporal logic task planning problems for mobile robots. We consider missions that require accomplishing multiple high-level sub-tasks, expressed in natural language (NL), in a temporal and logical order. To formally define the mission, we treat these sub-tasks as atomic predicates in a Linear Temporal Logic (LTL) formula. We refer to this task specification framework as LTL-NL. Our goal is to design plans, defined as sequences of robot actions, accomplishing LTL-NL tasks. This action planning problem cannot be solved directly by existing LTL planners due to the NL nature of atomic predicates. Therefore, we propose HERACLEs, a hierarchical neuro-symbolic planner that relies on a novel integration of (i) existing symbolic planners generating high-level task plans determining the order at which the NL sub-tasks should be accomplished; (ii) pre-trained Large Language Models (LLMs) to design sequences of robot actions for each sub-task in these task plans; and (iii) conformal prediction acting as a formal interface between (i) and (ii) and managing uncertainties due to LLM imperfections. We show, both theoretically and empirically, that HERACLEs can achieve user-defined mission success rates. We demonstrate the efficiency of HERACLEs through comparative numerical experiments against recent LLM-based planners as well as hardware experiments on mobile manipulation tasks. Finally, we present examples demonstrating that our approach enhances user-friendliness compared to conventional symbolic approaches.
@@ Abstract.
copyright: acmlicensedjournal: TCPSccs: Computing methodologies Robotic planningccs: Computing methodologies Planning under uncertaintyccs: Computer systems organization Robotic control
1. Introduction
@@
Designing autonomous agents with task planning capabilities is a long-standing goal in robotics and automation (Garrett et al., 2021; Antonyshyn et al., 2023). Achieving this goal requires the development of mission specification frameworks and planning algorithms capable of generating plans—sequences of robot actions—to accomplish missions. These frameworks should allow users to define missions unambiguously and in a user-friendly manner, while planners need to exhibit computational efficiency and be supported by correctness guarantees.
@@
Recently, several task planners have been proposed that can design correct-by-construction plans for complex missions with specified temporal and logical requirements using Linear Temporal Logic (LTL) (Sahin et al., 2019; Kantaros et al., 2022; Luo et al., 2021; Ulusoy et al., 2014; Sun and Shoukry, 2024; Kantaros and Zavlanos, 2020; Vasile and Belta, 2013; Ho et al., 2023; Zhou et al., 2023; Kantaros et al., 2020; Vasilopoulos et al., 2021; Chen et al., 2024; Leahy et al., 2022; Guo and Dimarogonas, 2015; Shoukry et al., 2017; Hashimoto et al., 2022; Kantaros and Zavlanos, 2017; Chen et al., 2018). However, this framework demands a significant amount of expertise and manual effort to rigorously define complex tasks. This is because defining LTL tasks requires specifying multiple atomic predicates (i.e., Boolean variables) to model desired low-level robot configurations and coupling them using temporal/Boolean operators. Additionally, the more complex the task requirements, the larger the number of predicates and temporal/logical operators needed to define the corresponding LTL formula. This not only compromises the user-friendliness but also increases the computational cost of designing robot plans (Kantaros and Zavlanos, 2020). On the other hand, Natural Language (NL) has emerged as a more user-friendly method for specifying robot missions. Early research in NL-based planning mainly focused on mapping NL to planning primitives (Tellex et al., 2011; Matuszek et al., 2010; Chen and Mooney, 2011; Howard et al., 2014) using statistical machine translation (Koehn, 2009) to identify data-driven patterns for translating free-form commands into a formal language defined by a grammar. However, these approaches were limited to structured state spaces and simple NL commands. Motivated by the remarkable generalization abilities of pre-trained Large Language Models (LLMs) across diverse task domains there has been increasing attention on utilizing LLMs for NL-based planning (Singh et al., 2023; Ren et al., 2023; Liang et al., 2023; Shah et al., 2023; Ding et al., 2023; Liu et al., 2023; Stepputtis et al., 2020; Li et al., 2022; Huang et al., 2022; Ahn et al., 2022; Ruan et al., 2023; Luo et al., 2023; Joublin et al., 2023; Yang et al., 2024b; Rana et al., 2023; Wang et al., [n. d.]). While LLM-based planning methods are more user-friendly than LTL-based planners, they have two main limitations: (i) they lack mission performance guarantees; and (ii) their ability to design correct plans deteriorates with increasing task complexity (Chen et al., 2023a). In this paper, we propose a new approach to mission specification and robot planning that combines the formal rigor of LTL-based planners with the user-friendliness of NL-based approaches.
@@
First, we discuss our proposed mission specification framework. Our specification framework, called LTL-NL, models complex tasks as collections of NL-described sub-tasks with logical and temporal dependencies expressed using standard Boolean and temporal operators. Each NL sub-task is treated as an atomic predicate—evaluated as true when successfully completed, and false otherwise. The key departure from conventional LTL-based methods lies in specifying predicates/sub-tasks using NL, at the level of intent, rather than low-level system states (see, e.g., (Schuppe and Tumova, 2020; Chen et al., 2012; Kantaros and Zavlanos, 2020; Vasile and Belta, 2013; He et al., 2015)). For instance, consider a task where a robot must first deliver a bottle of water to the kitchen table, and then deliver a Coke to the office desk. We can define two NL sub-tasks as predicates, ‘deliver a bottle of water to the kitchen table’ and ‘deliver a Coke to the office desk’, and use Boolean/temporal operators to express the logical and temporal relationship between them. In contrast, conventional LTL formulas require low-level predicates that are true if e.g., the robot is near the bottle or if it is holding the object. Our LTL-NL framework results in a more compact and user-friendly mission specification compared to conventional LTL formulas. In comparison to NL instructions, our framework enables the automatic decomposition of the overall mission (i.e., an LTL-NL specification) into multiple NL-based sub-tasks using existing symbolic planners (Kantaros et al., 2020; Vasilopoulos et al., 2021; Leahy et al., 2022). It is important to note that such a decomposition is highly challenging to perform in a correct-by-construction fashion for missions expressed exclusively using NL. This decomposition is critical for ensuring logical consistency and robust performance -particularly as mission complexity increases- enabling our planning framework to consistently outperform existing LLM-based planners.
Figure 1. This figure illustrates a robot operating in a post-disaster city. The high-level mission-”deliver either traffic cone 1, 2, or 3 to location A, and only then deliver fire hydrant 1 to location B”-is formulated with an LTL-NL formula $\phi=\diamond\pi_{4}\wedge(\neg\pi_{4}{\mathcal{U}}(\pi_{1}\vee\pi_{2}\vee\pi_{3}))$ with 4 NL subtasks ($\pi_{1}\sim\pi_{4}$). The symbolic planner initially generates the NL-subtask $\pi_{1}$ (‘deliver traffic cone 1 to location A’) along with the safety constraint $\neg\pi_{4}$ (‘do not move the fire hydrant 1 to location B’). As the robot executes the LLM-generated plan to accomplish this sub-task, it builds a grid map of the environment revealing that traffic cone 1 is blocked by surrounding cars, making it inaccessible. Consequently, the symbolic planner assigns a new sub-task, modeled by $\pi_{3}$ and $\neg\pi_{4}$, requiring the robot to deliver traffic cone 3 to A; see also Example 2.3 and the demonstrations in (Demonstrations, 2024).
@@
Second, we address the challenge of designing robot plans that complete LTL-NL tasks. Specifically, we consider a mobile robot equipped with various skills (e.g., mobility, manipulation, and sensing) tasked with missions expressed as LTL-NL specifications. Each predicate in the LTL-NL formula is satisfied if an NL-based sub-task is accomplished, requiring the robot to apply its skills to various semantic objects and regions of interest in the environment. For instance, the sub-task ‘deliver a bottle of water to the kitchen table’ requires the robot to move to the location where the bottle is, grasp it, move to the kitchen table, and release the object. Observe that a predicate in an LTL-NL formula is satisfied by a sequence of robot configurations/actions. While we assume that the locations and semantic labels of these objects are known, the geometric environmental structure is initially unknown. The latter may hinder the completion of certain NL-based sub-tasks (e.g., when an object is inaccessible); see Fig. 1. Our goal is to design robot plans, defined as a sequence of robot actions, satisfying the assigned LTL-NL missions.
@@
To address this planning problem, we propose a hierarchical neuro-symbolic planner consisting of (1) a symbolic planner, determining the next NL-encoded sub-task the robot should execute to make mission progress, and (2) a neural planner, which translates these sub-tasks into executable robot actions. In our implementation, we adopt the symbolic planner from (Kantaros et al., 2020; Vasilopoulos et al., 2021) due to its computational efficiency, though our framework remains compatible with other symbolic planners. Given the NL sub-task selected by the symbolic planner, the neural planner uses a pre-trained LLM to generate a robot plan that satisfies it (Touvron et al., 2023; Brown et al., 2020). It is possible that the initially unknown geometric structure of the environment may prevent the robot from fully executing a generated plan (e.g., due to obstacles blocking access to regions of interest). In such cases, a ‘help’ message is sent to the symbolic planner, which responds by issuing alternative sub-tasks to the neural planner, if available, (Kantaros et al., 2020; Vasilopoulos et al., 2021); see Fig. 1. Once a sub-task is completed, the symbolic planner selects the next one and forwards it to the neural planner for plan generation. This interaction between the symbolic and neural planners continues until the overall LTL-NL mission is completed. A key challenge in implementing the neural planner is that LLMs tend to hallucinate, i.e., to confidently generate incorrect outputs (LeCun, 2023). To reason about the correctness of the LLM-generated plans, inspired by (Ren et al., 2023; Kumar et al., 2023), we leverage conformal prediction (CP) (Angelopoulos and Bates, 2021), a statistical tool for uncertainty quantification in black-box models that has been recently applied to various safe autonomy tasks (Angelopoulos et al., 2023; Kumar et al., 2023; Ren et al., 2023; Mao et al., 2023; Muthali et al., 2023; Tonkens et al., 2023; Lindemann et al., 2023; Yang et al., 2024a; Qin et al., 2024). CP constructs on-the-fly prediction sets that contain the correct robot action with user-specified confidence. This allows the LLM to determine when it is uncertain about its predictions. In cases of high uncertainty, indicated by non-singleton prediction sets, the LLM seeks assistance from a user or the symbolic planner. This formal interface between the symbolic planner and the LLM, enabled by CP, gives rise to our planner, called HERACLEs, for HiERArchical ConformaL natural languagE planner. The generated plans are executed using existing low-level controllers. We show both theoretically and empirically that HERACLEs can achieve user-specified mission completion rates, thanks to CP, while requiring low help rates—significantly lower than related works (Ren et al., 2023), due to its neuro-symbolic architecture.
@@
Related works: (i) As discussed earlier, a substantial body of symbolic planners exists that can design plans satisfying LTL tasks (Sahin et al., 2019; Kantaros et al., 2022; Luo et al., 2021; Ulusoy et al., 2014; Sun and Shoukry, 2024; Kantaros and Zavlanos, 2020; Vasile and Belta, 2013; Ho et al., 2023; Zhou et al., 2023; Kantaros et al., 2020; Vasilopoulos et al., 2021; Chen et al., 2024; Leahy et al., 2022; Guo and Dimarogonas, 2015; Shoukry et al., 2017; Hashimoto et al., 2022; Kantaros and Zavlanos, 2017; Chen et al., 2018). We emphasize that these planners cannot solve the considered LTL-NL planning problem due to the NL nature of the predicates. Specifically, these works can design a sequence of NL predicates that need to be satisfied to accomplish the LTL-NL mission; however, they cannot design a sequence of robot actions (e.g., ‘grab the bottle of water at location A’, ‘go to the kitchen table’, ‘drop off the bottle’) that satisfies a NL-based predicate (e.g., ‘deliver a bottle of water to the kitchen table’), let alone an LTL-NL formula, as they cannot process NL. We note again that our approach is tightly integrated with these symbolic planners as they are used to generate a sequence of NL sub-tasks that should be accomplished to satisfy an LTL-NL formula. (ii) Related are also recent neural planners that can design plans for tasks, described solely in NL, using pre-trained LLMs; see e.g., (Singh et al., 2023; Ren et al., 2023; Liang et al., 2023; Shah et al., 2023; Ding et al., 2023; Liu et al., 2023; Stepputtis et al., 2020; Li et al., 2022; Huang et al., 2022; Ahn et al., 2022; Ruan et al., 2023; Luo et al., 2023; Joublin et al., 2023; Yang et al., 2024b; Rana et al., 2023; Wang et al., [n. d.]); a recent survey can be found in (Pallagani et al., 2024; Hunt et al., 2024). However, the ability of LLMs to design correct plans drops significantly as the mission complexity increases as shown in (Chen et al., 2023a; Valmeekam et al., 2024) as well as in our experiments. Additionally, as discussed earlier, the majority of these planners lack correctness and mission completion guarantees. Among related works, the most closely aligned is KnowNo (Ren et al., 2023), which also employs CP to quantify uncertainty in LLM outputs and enables robots to seek help from users to achieve desired mission completion rates. However, as demonstrated in our comparative experiments, HERACLEs achieves significantly lower help rates, primarily due to its task decomposition enabled by the symbolic planner. Moreover, HERACLEs can identify specific sub-tasks that cannot be completed due to the geometric constraints of the environment and adapt by requesting alternative sub-tasks from the symbolic planner. This reactive capability is not supported by (Ren et al., 2023), and as a result, it cannot readily adapt to unknown environments. (iii) Neuro-symbolic planners that integrate LLMs with symbolic planners have also been proposed recently in (Chen et al., 2023a; Quartey et al., 2024; Dagan et al., 2023; Fuggitti and Chakraborti, 2023; Chen et al., 2023c; Cosler et al., 2023; Dai et al., 2023). Their key idea is to employ LLMs to translate NL commands into temporal logic specifications. Once this translation is completed, existing symbolic planners can be employed to design correct plans. A key limitation is that the translation process is not supported by correctness guarantees. As a result, the resulting plan may not satisfy the original NL instructions. In contrast, HERACLEs employs LLMs and symbolic planners in a fundamentally different way while CP allows the design of plans that are probabilistically correct and, therefore, they can achieve desired mission completion rates.
@@
Comparative Evaluations: We provide extensive comparative experiments to demonstrate the benefits of the proposed method. First, we compare HERACLEs against KnowNo (Ren et al., 2023) as it is the closest work to ours. As discussed earlier, KnowNo can also achieve desired mission success rates by incorporating CP in the plan synthesis phase and allowing robots to ask for help in cases of high uncertainty. Since KnowNo requires the mission to be expressed in NL, as opposed to LTL-NL, in our comparisons, we manually translate LTL-NL missions into NL instructions serving as inputs to (Ren et al., 2023). Our empirical analysis demonstrates that HERACLEs achieves significantly lower help rates than (Ren et al., 2023) to attain the same mission success rate. We attribute the performance gap between KnowNo and HERACLEs primarily to the LTL-NL framework and the symbolic planner decomposing missions into multiple smaller/simpler NL-based sub-tasks that LLMs can handle more efficiently. Second, we provide examples demonstrating the user-friendliness of LTL-NL formulas compared to corresponding LTL specifications. Specifically, we show that LTL-NL formulas require significantly fewer predicates than LTL formulas defined over low-level configuration-based predicates modeling the same task. This difference in the ‘length’ of LTL and LTL-NL formulas becomes more pronounced as the complexity of the mission requirements increases. Here we would like to note that direct comparisons with other temporal logic planners are not possible, as they cannot handle LTL-NL formulas, as discussed earlier. Similarly, comparisons with the neuro-symbolic approaches mentioned above are not straightforward, since they address NL-to-LTL translation problems, while our focus is on solving planning problems once an LTL-NL formula is specified. In Appendix A, we also provide additional comparisons against the LLM-based planners proposed in (Singh et al., 2023; Ahn et al., 2022; Chen et al., 2023b), which, however, do not allow robots to ask for help and, therefore, they cannot achieve desired mission completion rates. These comparative experiments show that the ability of these approaches to design correct plans tends to decrease as the number of temporal and logical requirements in the mission increases. To the contrary, the planning performance of HERACLEs consistently outperforms these baselines (even when the help mode is deactivated).
@@
Contribution: The contribution of the paper can be summarized as follows. First, we propose a new mission specification framework, called LTL-NL, to define complex high-level robot tasks. The advantage of LTL-NL over conventional LTL is its user-friendliness, as it requires fewer predicates to define the same task while the predicates can be defined using NL rather than low-level system configurations. Second, we introduce HERACLEs, a new neuro-symbolic planner that integrates symbolic planners with LLMs to design plans satisfying missions encoded as LTL-NL formulas in partially known environments. Third, we show, both theoretically and empirically, that HERACLEs can achieve user-specified mission success rates through a formal interface, employing CP, between the LLM and the symbolic planner. Fourth, we conduct extensive comparative experiments showing that HERACLEs outperforms state-of-the-art LLM-based planners both in generating correct plans and in minimizing human intervention (help rates). Finally, we provide hardware experiments on mobile manipulation tasks to demonstrate the effectiveness of our method.
2. Problem Formulation